OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 2x – 7y = 1 x – 8y = –4

OpenStudy (turingtest):

third time I've seen this problem today lol from (2) x=8y-4 sub into (1) 2(8y-4)-7y=1 16y-8-7y=1 9y=9 y=1 sub this answer back into (2) x-8(1)=-4 x=4

OpenStudy (anonymous):

you guys should get an account for the entire class, saves us work

OpenStudy (turingtest):

I know, right?